OpenStudy (anonymous):

Find the volume of a prism if the base is a square with sides measuring 7 and an atitude of 3

OpenStudy (anonymous):

V=7*7*3 v= 147

OpenStudy (anonymous):

find the area of the base of a prism if the volume is 42 and the height is 3 and thank you very much

OpenStudy (anonymous):

A = L*W V= A*H 42= A*3 42/3=A A=14 If I deserve medal then give me one

OpenStudy (anonymous):

lol a right prism has a rhombus for its base. if the diagonals of the rhombus are 12 ft and 16ft and the height of the prism is 20 ft what is the total surface area and volume of this prism

OpenStudy (anonymous):

new qustion?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

be back